Source of Variation SS df MS F P-value F crit
Rows 4.80 4 c f j h
Columns a b d g k i
Error 15.09 12 e
Total 114.07 19

In this Two Way ANOVA, the conclusion you arrive at is that mean blocks are not all the same.

True

False
Answer #1

Null hypothesis: no difference in block means

Then F statistic is g, ( blocks are columns) & follow F distribution with degree of freedom: b, 12

Now b is 3

As b = 19-12-4, g= d/e

d is a/b, e is 15.09/12 = 1.2575

a= 114.07-15.09-4.80 = 94.18

Thus d= 94.18/3 = 31.3934

Thus F = g = 31.3934/1.2575

F=24.964

Thus F statistic is highly significant, hence p value is very small, thus null hypothesis is rejected.

Answer is True